OpenStudy (anonymous):

I know how to find the derivative from sin(x) to cos(x) of sqrt(4-x) dx by splitting the integral at 0 and rewriting it and using the integral properties. What I am having trouble seeing is what is a graphical interpretation of an integral like that? Can you point me towards a resource?

OpenStudy (anonymous):

\[\int\limits_{\sin x}^{\cos x} \sqrt(4-x) dx\] This kind of integral

OpenStudy (anonymous):

The integrand doesn't really matter